# Machine-generated text detection prevents language model collapse

### v5

@@ Appendix C Informed sampling of training data
SmolLM 135M 10 10
SmolLM 360M 5 10
SmolLM 1.7B 3 1
@@
Figure S1: Perplexity and accuracy over generations 00 to $9$ of fully synthetic recursive training for varying mixing coefficients ($\alpha$, $\beta$, $\gamma$) using top-$k$ decoding.
@@
Figure S2: Classification score distribution of the machine-generated text detector on the dataset from generation 00.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure S3: Perplexity distribution for machine-generated text of SmolLM2 at generations 00, $1$, and $9$ in the partially synthetic scenario $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$. Perplexity is evaluated using the model trained on human text ($p^{0}$).
@@
Figure S4: GPT-2 (top) and SmolLM2 (bottom) under partially synthetic recursive training $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$ for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
$1.7$B Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$360$M Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$135$M Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure S5: SmolLM2 model size variants ($1.7$B, $360$M, $135$M) under partially synthetic recursive training $\left(\alpha\!=\!1,\beta\!=\!1,\gamma\!=\!0\right)$ for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
@@
Figure S6: GPT-2 (top) and SmolLM2 (bottom) under partially synthetic recursive training with cross-generational data $\left(\alpha\!=\!0.5,\beta\!=\!0.5,\gamma\!=\!0.5\right)$ for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
@@
Gen 0 Gen 9 Gen 0 Gen 9 Gen 0 Gen 9 Gen 0 Gen 9 Gen 0 Gen 9 Gen 0 Gen 9
GPT-2 baseline top-$k$5, 1, 0 $29.23$ $31.85$ $38.80$ $37.56$ $84.46$ $82.16$ $39.10$ $39.55$ $93.31$ $91.10$ $51.30$ $46.16$
.5, .5, .5 $29.22$ $31.33$ $38.83$ $37.84$ $85.66$ $84.35$ $38.94$ $40.64$ $94.07$ $92.00$ $51.06$ $45.38$
1, 1, 0 $29.25$ $29.92$ $38.78$ $38.34$ $84.66$ $82.68$ $39.19$ $40.20$ $93.24$ $92.69$ $50.99$ $48.09$
sampling .5, 1, 0 $29.25$ $32.27$ $38.78$ $38.01$ $94.86$ $97.77$ $24.09$ $14.52$ $91.08$ $46.87$ $40.64$ $23.69$
.5, .5, .5 $29.26$ $31.82$ $38.77$ $38.03$ $94.95$ $97.50$ $24.15$ $17.02$ $90.07$ $76.01$ $41.02$ $29.08$
1, 1, 0 $29.25$ $38.79$ $38.79$ $38.63$ $94.88$ $96.86$ $24.04$ $18.12$ $91.18$ $75.64$ $40.71$ $31.49$
ours top-$k$5, 1, 0 $29.25$ $29.53$ $38.78$ $38.45$ $84.43$ $85.89$ $39.43$ $38.85$ $94.59$ $91.52$ $50.79$ $50.33$
.5, .5, .5 $29.25$ $29.47$ $38.77$ $38.49$ $85.18$ $86.40$ $39.28$ $38.90$ $94.75$ $95.67$ $51.31$ $49.99$
1, 1, 0 $29.25$ $28.59$ $38.77$ $38.91$ $84.81$ $85.65$ $38.85$ $38.76$ $94.92$ $93.95$ $51.28$ $51.34$
pure sampling .5, 1, 0 $29.25$ $29.88$ $38.79$ $38.58$ $94.87$ $96.50$ $24.30$ $19.85$ $92.92$ $81.58$ $40.68$ $35.59$
.5, .5, .5 $29.26$ $29.74$ $38.76$ $38.60$ $94.92$ $96.46$ $24.09$ $20.66$ $91.16$ $88.46$ $40.98$ $36.48$
1, 1, 0 $29.24$ $28.84$ $38.78$ $39.00$ $94.61$ $96.17$ $24.01$ $21.87$ $91.76$ $88.06$ $40.78$ $38.10$
SmolLM2 $350$M baseline top-$k$5, 1, 0 $13.96$ $14.69$ $47.58$ $47.30$ $82.59$ $73.95$ $52.68$ $54.53$ $91.58$ $78.42$ $55.31$ $58.10$
.5, .5, .5 $13.96$ $14.64$ $47.59$ $47.30$ $82.76$ $76.54$ $52.57$ $53.97$ $91.03$ $86.15$ $55.38$ $57.84$
1, 1, 0 $13.96$ $14.10$ $47.57$ $47.57$ $82.51$ $76.90$ $52.47$ $54.17$ $91.32$ $81.11$ $55.57$ $57.24$
temperature .5, 1, 0 $13.96$ $14.69$ $47.59$ $47.37$ $83.16$ $66.16$ $51.21$ $54.46$ $83.67$ $70.15$ $52.59$ $57.81$
.5, .5, .5 $13.96$ $14.60$ $47.58$ $47.36$ $83.74$ $72.03$ $51.22$ $54.15$ $86.94$ $81.27$ $52.80$ $56.86$
1, 1, 0 $13.96$ $14.06$ $47.58$ $47.64$ $83.31$ $73.11$ $51.16$ $53.96$ $84.55$ $76.27$ $52.74$ $55.85$
ours top-$k$5, 1, 0 $13.96$ $14.02$ $47.58$ $47.62$ $82.75$ $79.17$ $52.58$ $53.89$ $89.33$ $90.39$ $55.53$ $57.67$
.5, .5, .5 $13.96$ $14.00$ $47.59$ $47.59$ $82.52$ $79.79$ $52.64$ $53.74$ $88.69$ $87.92$ $55.39$ $57.41$
1, 1, 0 $13.96$ $13.67$ $47.60$ $47.83$ $82.93$ $81.67$ $52.74$ $53.23$ $85.09$ $89.72$ $55.53$ $56.90$
temperature .5, 1, 0 $13.96$ $14.11$ $47.58$ $47.63$ $83.45$ $75.84$ $51.15$ $53.67$ $87.80$ $84.82$ $53.03$ $56.74$
.5, .5, .5 $13.96$ $14.13$ $47.58$ $47.58$ $82.99$ $77.39$ $51.12$ $53.08$ $86.20$ $81.00$ $52.98$ $56.09$
1, 1, 0 $13.96$ $13.75$ $47.57$ $47.83$ $83.43$ $79.36$ $51.21$ $52.99$ $88.16$ $86.94$ $52.94$ $55.66$
SmolLM2 $135$M baseline top-$k$ 1, 1, 0 $19.02$ $19.35$ $44.08$ $44.04$ $79.08$ $70.99$ $53.08$ $54.87$ $85.18$ $72.73$ $57.34$ $59.49$
@@
Table S2: Test performance (perplexity and accuracy) and data quality at generation 00 and generation $9$. Results are shown for top-$k$ decoding and pure sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ ($\uparrow$ / $\downarrow$: higher / lower is better).

### v6

@@ Appendix C Informed sampling of training data
SmolLM 135135M $10$ $10$
SmolLM 360360M 55 $10$
SmolLM 1.71.7B 33 11
@@
Figure S1: Perplexity and accuracy over generations 0 to 99 of fully synthetic recursive training for varying mixing coefficients $\left(\alpha,\beta,\gamma\right)$ using top-$k$ decoding.
@@
Figure S2: Classification score distribution of the machine-generated text detector on the dataset from generation 0.
@@
Figure S3: GPT-2 (top) and SmolLM2 (bottom) under partially synthetic recursive training $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$ for 10 generations. The baseline is equivalent to training on all the data in the pool and the ‘Oracle’ performance represents a perfect AI text detector that filters all synthetic samples.
@@
Figure S4: GPT-2 (top) and SmolLM2 (bottom) under partially synthetic recursive training with cross-generational data $\left(\alpha\!=\!0.5,\beta\!=\!0.5,\gamma\!=\!0.5\right)$ for 10 generations. The baseline is equivalent to training on all the data in the pool and the ‘Oracle’ performance represents a perfect AI text detector that filters all synthetic samples.
1.71.7B Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
360360M Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
135135M Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure S5: SmolLM2 model size variants (1.71.7B, 360360M, 135135M) under partially synthetic recursive training $\left(\alpha\!=\!1,\beta\!=\!1,\gamma\!=\!0\right)$ for 10 generations. The baseline is equivalent to training on all the data in the pool and the ‘Oracle’ performance represents a perfect AI text detector that filters all synthetic samples.
@@ 1.71.7B Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Gen 0 Gen 99 Gen 0 Gen 99 Gen 0 Gen 99 Gen 0 Gen 99 Gen 0 Gen 99 Gen 0 Gen 99
GPT-2 baseline top-$k$ .5.5, 11, 0 $29.23$ $31.85$ $38.80$ $37.56$ $84.46$ $82.16$ $39.10$ $39.55$ $93.31$ $91.10$ $51.30$ $46.16$
.5.5, .5.5, .5.5 $29.22$ $31.33$ $38.83$ $37.84$ $85.66$ $84.35$ $38.94$ $40.64$ $94.07$ $92.00$ $51.06$ $45.38$
11, 11, 0 $29.25$ $29.92$ $38.78$ $38.34$ $84.66$ $82.68$ $39.19$ $40.20$ $93.24$ $92.69$ $50.99$ $48.09$
pure sampling .5.5, 11, 0 $29.25$ $32.27$ $38.78$ $38.01$ $94.86$ $97.77$ $24.09$ $14.52$ $91.08$ $46.87$ $40.64$ $23.69$
.5.5, .5.5, .5.5 $29.26$ $31.82$ $38.77$ $38.03$ $94.95$ $97.50$ $24.15$ $17.02$ $90.07$ $76.01$ $41.02$ $29.08$
11, 11, 0 $29.25$ $38.79$ $38.79$ $38.63$ $94.88$ $96.86$ $24.04$ $18.12$ $91.18$ $75.64$ $40.71$ $31.49$
ours top-$k$ .5.5, 11, 0 $29.25$ $29.53$ $38.78$ $38.45$ $84.43$ $85.89$ $39.43$ $38.85$ $94.59$ $91.52$ $50.79$ $50.33$
.5.5, .5.5, .5.5 $29.25$ $29.47$ $38.77$ $38.49$ $85.18$ $86.40$ $39.28$ $38.90$ $94.75$ $95.67$ $51.31$ $49.99$
11, 11, 0 $29.25$ $28.59$ $38.77$ $38.91$ $84.81$ $85.65$ $38.85$ $38.76$ $94.92$ $93.95$ $51.28$ $51.34$
pure sampling .5.5, 11, 0 $29.25$ $29.88$ $38.79$ $38.58$ $94.87$ $96.50$ $24.30$ $19.85$ $92.92$ $81.58$ $40.68$ $35.59$
.5.5, .5.5, .5.5 $29.26$ $29.74$ $38.76$ $38.60$ $94.92$ $96.46$ $24.09$ $20.66$ $91.16$ $88.46$ $40.98$ $36.48$
11, 11, 0 $29.24$ $28.84$ $38.78$ $39.00$ $94.61$ $96.17$ $24.01$ $21.87$ $91.76$ $88.06$ $40.78$ $38.10$
SmolLM2 350350M baseline top-$k$ .5.5, 11, 0 $13.96$ $14.69$ $47.58$ $47.30$ $82.59$ $73.95$ $52.68$ $54.53$ $91.58$ $78.42$ $55.31$ $58.10$
.5.5, .5.5, .5.5 $13.96$ $14.64$ $47.59$ $47.30$ $82.76$ $76.54$ $52.57$ $53.97$ $91.03$ $86.15$ $55.38$ $57.84$
11, 11, 0 $13.96$ $14.10$ $47.57$ $47.57$ $82.51$ $76.90$ $52.47$ $54.17$ $91.32$ $81.11$ $55.57$ $57.24$
temperature .5.5, 11, 0 $13.96$ $14.69$ $47.59$ $47.37$ $83.16$ $66.16$ $51.21$ $54.46$ $83.67$ $70.15$ $52.59$ $57.81$
.5.5, .5.5, .5.5 $13.96$ $14.60$ $47.58$ $47.36$ $83.74$ $72.03$ $51.22$ $54.15$ $86.94$ $81.27$ $52.80$ $56.86$
11, 11, 0 $13.96$ $14.06$ $47.58$ $47.64$ $83.31$ $73.11$ $51.16$ $53.96$ $84.55$ $76.27$ $52.74$ $55.85$
ours top-$k$ .5.5, 11, 0 $13.96$ $14.02$ $47.58$ $47.62$ $82.75$ $79.17$ $52.58$ $53.89$ $89.33$ $90.39$ $55.53$ $57.67$
.5.5, .5.5, .5.5 $13.96$ $14.00$ $47.59$ $47.59$ $82.52$ $79.79$ $52.64$ $53.74$ $88.69$ $87.92$ $55.39$ $57.41$
11, 11, 0 $13.96$ $13.67$ $47.60$ $47.83$ $82.93$ $81.67$ $52.74$ $53.23$ $85.09$ $89.72$ $55.53$ $56.90$
temperature .5.5, 11, 0 $13.96$ $14.11$ $47.58$ $47.63$ $83.45$ $75.84$ $51.15$ $53.67$ $87.80$ $84.82$ $53.03$ $56.74$
.5.5, .5.5, .5.5 $13.96$ $14.13$ $47.58$ $47.58$ $82.99$ $77.39$ $51.12$ $53.08$ $86.20$ $81.00$ $52.98$ $56.09$
11, 11, 0 $13.96$ $13.75$ $47.57$ $47.83$ $83.43$ $79.36$ $51.21$ $52.99$ $88.16$ $86.94$ $52.94$ $55.66$
SmolLM2 135135M baseline top-$k$ 11, 11, 0 $19.02$ $19.35$ $44.08$ $44.04$ $79.08$ $70.99$ $53.08$ $54.87$ $85.18$ $72.73$ $57.34$ $59.49$
@@
Table S2: Test performance (perplexity and accuracy) and data quality at generation 0 and generation $9$. Results are shown for top-$k$ decoding and pure sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ ($\uparrow$ / $\downarrow$: higher / lower is better).